You want to find the standard deviation of the internal body temperatures of a group of healthy adults. Here are the temperatures, in degrees Fahrenheit:

99.2   98.9   98.2   97.3   98.9   98.4   97.2   99.1

Find the mean of this data set: Answer?

Now, use the mean to find the deviation for each score. The deviation is the score minus the mean.

Data Value Deviation = Data Value - Mean
99.2 Answer?
98.9 Answer?
98.2 Answer?
97.3 Answer?
98.9 Answer?
98.4 Answer?
97.2 Answer?
99.1 Answer?

What is the sum of all the deviations? Answer?
Answer #1

Mean=(99.2+ 98.9+98.2+97.3+98.9 +98.4+97.2+99.1)/8

Mean=787.2/8

mean of this data set=98.4

Data value Deviation=Data value -mean Deviation ^2
99.2 99.2-98.4=0.8 0.64
98.9 98.9-98.4=0.5 0.25
98.2 98.2-98.4=-0.2 0.04
97.3 97.3-98.4=-1.1 1.21
98.9 98.9-98.4=0.5 0.25
98.4 98.4-98.4=0 0
97.2 97.2-98.4=-1.2 1.44
99.1 99.1-98.4=0.7 0.49
Sum=0 4.32

Sum of deviation=0.8+0.5-0.2-1.1+0.5+0-1.2+0.7=0 d

Population std deviation deviation=Sqrt(sum(Data value-mean)2/n)=Sqrt(4.32/8)=0.734847

Sample deviation deviation=Sqrt(sum(Data value-mean)2/n-1) =sqrt(4.32/7)=0.7855